The Burdell Company is a small manufacturing company that uses gear assemblies to produce four different models of mountain bikes. One of these gear assemblies, the "Smooth Shifter", is used for the two most expensive of Burdell's four models, and has an estimated weekly demand of 7 units. Currently, this assembly is purchased in lots of 100 units. Burdell estimates the cost to place an order is $40, and the holding cost for each assembly is $5/unit/month. The company operates 300 days per year. (Assume 1 year = 12 months = 52 weeks)

a. What is the total annual cycle-inventory cost based on current ordering policy?

Answers to A and B are below

b. What is the economic order quantity for this assembly?

c.What is the total annual cycle-inventory cost based on EOQ policy?

d. What is the time between orders measured in days, if EOQ policy is used?
Answer #1

a:

Annual demand, D= 7*52 weeks 364
Order cost or setup cost, S $40
holding cost per year, H= $60.00
Order quantity, Q= 100
annual holding cost= Q*H/2 $3,000.0

the total annual cycle-inventory cost based on current ordering policy= $3000

demand per week 7
Days per year= 300
Annual demand, D= 7*52 364
Order cost or setup cost, S $40
holding cost per year, H= $60.00
Answer b Economic Order quantity, Q= squareroot(2*S*D/H) 22
#orders per year = D/Q 16.52
Answer c annual holding cost= Q*H/2 $660.9
Answer d order cycle= days per year/orders per year 18.16